The electric potential in a region of space is V=( 100 x2− 240 y2)V, where x and y are in meters. What is the direction of the electric field at (x,y)=(1.0m,3.0m)? Give the direction as an angle (in degrees) counterclockwise from the positive x-axis.
Answer #1

V = 100x^2 - 240y^2

Ex = -dV/dx

= -200x

Ey = -dV/dy

= 480y

E = Ex i + Eyj

= -200x i + 480y j

E(1,3) = -200*1 i + 480*3 j

- -200i + 1440 j

Direction = 180 - tan^-1 ( 1440/200) = 97.9o from the x-axis in counter clockwise direction